### v1

# CGCOD: Class-Guided Camouflaged Object Detection

Abstract
Camouflaged Object Detection (COD) is designed to identify objects that blend seamlessly with their surroundings. Due to the complexity of camouflaged objects (such as shape, color, and texture), their semantic cues are often blurred or completely lost, posing a significant challenge for COD. Existing COD methods often rely on visual features, which are not stable enough in changeable camouflage environments. This instability leads to false positives and false negatives, resulting in incomplete or inaccurate segmentation results. In this paper, to solve this problem, we propose a new task, Class-Guided Camouflaged Object Detection (CGCOD), which extends the traditional COD task by introducing object class knowledge, significantly improving the robustness and segmentation accuracy of the model in complex environments. Toward this end, we construct a dataset, CamoClass, containing the camouflaged objects in the real scenes and their corresponding class annotation. Based on this, we propose a multi-stage framework CGNet which consists of a plug-and-play class prompt generator and a class-guided detector. Under the guidance of textual information, CGNet enables efficient segmentation. It is worth emphasizing that for the first time, we extend the object class annotations on existing COD benchmark datasets, and introduce a flexible framework to improve the performance of the existing COD model under text guidance.
1 Introduction
Accurate object perception in visual scenes is essential for image-level semantic reasoning and cognition. However, camouflaged objects often blend seamlessly into their surroundings, obscuring their appearance or shape, which poses a significant challenge for object perception. This makes capturing the semantic features of camouflaged objects particularly challenging. Addressing this issue holds significant value in domains such as medical image segmentation Fan et al. (2020b), smart agriculture Rustia et al. (2020), and environmental protection Nafus et al. (2015).
Figure 1: Visualization of the proposed CGNet illustrating its segmentation performance guided by various textual information.
In the early stages of COD, algorithms primarily rely on hand-crafted features to perform the detection process Galun et al. (2003), which exhibit limited generalization capabilities. With the rise of deep learning, existing COD methods have significant advanced by incorporating prior information such as edge Sun et al. (2022), gradient Zhang et al. (2023), and uncertainty Lyu et al. (2024), or employing multi-task learning He et al. (2023) to capture rich semantics. Despite substantial advancements across various scenes, these methods still face significant challenges, including limited robustness, low detection accuracy, and inadequate generalization capability. We argue that these challenges can be attribute to two aspects: (1) the inherently inconspicuous nature of camouflaged objects, which allows them to blend seamlessly into their surroundings, and (2) the complexity and variability of the environment conditions, such as occlusion and background saliency, which further hinder accurate detection.
These challenges arise from relying solely on purely visual paradigm, which fail to capture the full semantic characteristics of camouflaged objects. This limitation becomes particularly pronounced in scenes where objects closely resemble their backgrounds or are heavily occluded. Camouflaged objects exhibit strong visual similarity to their background (e.g., shape, color, scale), yet maintain significant semantic differences However, existing benchmark datasets primarily focus on visual information and lack stable, explicit features to effectively emphasize these semantic differences. As a result, purely visual models often suffer from high rates of false positives (e.g., misclassifying background regions as objects) and false negatives (e.g., failing to detect actual objects), leading to a substantial decline in detection accuracy. Moreover, various factors such as environmental lighting, interference, and occlusion, can profoundly affect the visual appearance of the object, making it challenging for models to consistently extract distinctive object features. For instance, partial occlusion distorts the original structure and contour of the object, reducing the model’s ability to recover complete semantic information from limited visual cues.
In recent years, integrating visual and textual features has proven effective for enhancing image content understanding, thereby improving performance in various multi-modal tasks Tan and Bansal (2019); Lu et al. (2019). Targeted textual prompts provide explicit semantic cues that help models focus on object-related features and reduce background interference Radford et al. (2021). However, existing COD benchmark datasets provide only visual information, while current visual-textual COD methods Zhang et al. (2024) depend on coarse-grained text generated by large language models. Therefore, it is of great significance for the COD community to build a dataset that integrates both visual and textual information and develop a new COD framework.
Motivated by the above discussion, we introduce a new task, Class-Guided Camouflaged Object Detection (CGCOD). This task accurately focus on relevant regions that likely contain camouflaged objects by leveraging the specific class text, thereby improving model’s detection accuracy and robustness. To support our research, we construct a new dataset, CamoClass. It integrates multiple COD datasets, including CAMO Le et al. (2019), COD10K Fan et al. (2022), CHAMELEON Przemysław et al. (2018), and NC4K Lv et al. (2021). We annotate all images with their corresponding class labels. To further explore the role of class guidance in COD, we propose a new method CGNet, which consist a plug-and-play class prompt generator (CPG) and a novel high-performance baseline class-guided detector (CGD). The CGD is designed to seamlessly integrate with CPG. In particular, we introduce a Class Semantic Guidance (CSG) consisting of two multi-head self-attention (MHSA Vaswani (2017)) modules as a bridge between CPG and CGD. The CSG aims to fuse information from multiple sources and reduce domain differences between them. Fig. 1 shows the ability of CGNet for the CGCOD task. In the CPG, we introduce a cross-modal multi-head attention module to enhance class semantics by facilitating interactions between visual and textual features. Additionally, our multi-level visual collaboration module (MVCM) ensures comprehensive alignment and integration of these features at various depths, enabling cooperative extraction of object semantics. Together, these modules generate class prompt features that guide the COD detector for precise segmentation. In the CGD, we design a semantic consistency module (SCM) to reduce false negatives and positives by leveraging the class prompt features from the CPG. CGNet further explores multi-modal feature correlations, suppresses background noise, and delivers accurate predictions.
The main contributions of our paper are as follows:
  • We propose a new task, CGCOD, which is the first time to fully explore the performance of COD model under class guidance and provide a new insight for traditional COD approaches.
  • We construct a novel dataset CamoClass by integrating existing COD datasets and annotate the class description of the camouflaged object.
  • We design CGNet, a new method featuring a plug-and-play class prompt generator that enhances existing COD models, and a high-performance baseline class-guided detector optimized for the CPG.
  • Comprehensive experiments demonstrate the superior performance of CGNet, which not only outperforms existing state-of-the-art methods, but also incorporates the CPG plug-and-play module to enhance the performance of other detectors.
2 Related Works
Referring Image Segmentation (RIS). RIS is a fundamental computer vision task aiming to segment object regions in an image based on natural language descriptions. Early RIS methods are often built upon CNN-LSTM architectures Hu et al. (2016). They encode image and text information separately, and leverage simple attention mechanisms. This limits cross-modal interaction, making it challenging for models to selectively focus on object-relevant features, resulting in coarse segmentation. For example, Wang et al. (2022) introduces a vision-language decoder that maps fine-grained text semantics to pixel-level representations for improved modality alignment. Tang et al. (2023) employs a mask classification framework with token-based querying to enhance object-level understanding and cross-modal reasoning. Yue et al. (2024) uses a word-guided dual-branch structure to dynamically align visual features with linguistic cues, boosting RIS performance. Although our method also integrates textual and visual modalities, our objective differs substantially from traditional RIS. RIS tasks use natural language to segment clearly described objects, whereas our task combines object class knowledge to detect camoflaged objects blending into the background, making it more challenging. This fundamental difference in goals and application distinguishes our work from RIS research.
Camouflaged Object Detection (COD). In nature, animals use camouflage as a defense mechanism to evade predators Copeland and Trivedi (1997). The high complexity of camouflaged strategies makes COD a consistently challenging task. Early COD methods relied on handcrafted features like texture Galun et al. (2003), gradient Zhang et al. (2016), and color Pulla Rao et al. (2020) to distinguish foreground from background but suffered from limited feature extraction capabilities, especially in complex scenes. In contrast, deep learning-based approaches, such as CNNs and Transformers, automatically learn camouflaged object features, offering significant improvements. For example, Fan et al. (2020a) uses densely connected features and receptive fields to mimic a predator’s search and recognition process. Pang et al. (2022) integrates multi-scale features and hierarchical units to emulate human visual patterns when observing blurry images, capturing hybrid-scale semantics. Sun et al. (2022) leverages edge cues to guide the model in perceiving object structures. Song et al. (2023) introduces a non-local token enhancement mechanism and a feature compression decoder to enhance feature interactions. He et al. (2023) employs a learnable wavelet transform to decompose features into different frequency domains, effectively distinguishing foreground from background. With the emergence of large models, Yu et al. (2024) utilizes SAM’s zero-shot capability by combining depth and RGB features for precise segmentation. Additionally, Zhang et al. (2024); Ruan et al. (2024) use environment descriptions generated by large language models to investigate how background camouflage attributes affect object detection. These approaches demonstrate deep learning’s flexibility and innovation in the COD task, offering diverse solutions.
Despite many advances, existing COD methods rely too heavily on visual features, which fail to adequately address the complex characteristics of camouflaged objects. This limitation may not only lead to poor model robustness, but also cause misrecognition problems.
In contrast, our work introduces class-level textual information into the COD domain, along with a new benchmark dataset.
By explicitly incorporating class semantics, our model can more accurately focus on the object region in the image and effectively reduce the interference of the complex environment on object perception.
Figure 2: Class distribution in the proposed CamoClass dataset.
3 Proposed CamoClass Dataset
In order to advance the study of visual-textual approaches in COD, we propose a new dataset, CamoClass. The focus of this research is to improve object detection in camouflage scenes by integrating the category knowledge of objects. This method extends the traditional idea of COD and provides a new benchmark. It should be emphasized that our innovation lies not only in the class annotation itself, but also in how to use class annotation to provide COD with new perspective and idea in visual-textual interaction without changing the existing model training paradigm.
Annotation Strategy. CamoClass contains 8,663 samples from 128 class, of which 44 are multi-class samples. Although the number of images of each class is uneven (as shown in Fig. 2), reflecting the varying frequencies of different camouflaged objects in the real world, this also poses challenges for the model’s generalization and robustness. Each sample consists of an image ($\mathbf{I}$), ground truth ($\mathbf{GT}$), and edge information ($\mathbf{E}$). In the process of labeling, we adopt strict class definition and standardization strategy to reduce ambiguity and ensure clarity and consistency of class definition, and unify the treatment of fine-grained class such as “small fish” and “black cat”. Grouping them into broader class (such as “fish” and “cat”) simplifies the class space and makes it easier to identify object. For images containing multiple-class objects, we provide an independent segmentation mask for each class. For example, an image containing both a "dog" and a "cat" will have separate masks for each class, ensuring accurate and high-quality annotation.
Flexible Application. CamoClass is highly flexible and extensible. Researchers can continue to use the training and test partitioning of the original COD dataset to maintain comparability with previous studies and ensure that the introduction of new information does not affect existing experimental paradigms. In addition, we integrate these sub-datasets to form a large COD dataset with diverse classes, and further divide the test set into seen classes $C_{seen}$ (present in the training set) and unseen classes $C_{unseen}$ (not present in the training set). The integrated dataset provides sufficient training data for the class-guided COD task, supporting the joint semantic modeling of text and image, and the location of camouflaged object in complex scenes. It is expected that this large-scale dataset will become a cornerstone of community research, driving the development of the COD field in open-vocabulary Pang et al. (2025), referring Ji et al. (2024), and zero-shot learning Ren et al. (2023).
Figure 3: Overview of the proposed CGNet.
4 Proposed Method
When students are in class, they don’t just rely on memorization, but externalize information by recording key concepts or points of knowledge. This focus on clearly targeted content helps them to grasp the learning points in a targeted way. Inspired by this principle, we introduce the CGNet to alleviate the limitations faced by the purely visual COD model. Fig. 3 provides an overview of our network architecture. In particular, the input of the CGNet is composed of tow parts, which include image and class-level textual information of the camouflaged object, termed as $I_{r}$ and $I_{t}$. They also comes from the dataset CamoClass, which are fed into the class prompt generator (CPG) and class-guided detector (CGD) to generate class-prompt feature $G_{c}$ and the segmentation result $P_{1}$.
4.1 Class Prompt Generator (CPG)
Overview. The CPG is designed to integrate both visual and textual information to generate prompt features for object classes. By leveraging textual information, it enhances the semantic understanding of the object’s context in the image, which helps mitigate the impact of visually similar but irrelevant objects. This mechanism significantly reduces both false positives and false negatives, thus improving the precision of the COD task. Specifically, in this phase, we first employ a frozen CLIP textual encoder to generate textual embeddings $F_{t}$ from the class-level textual information. Simultaneously, we extract multi-level visual features $\{F_{i}\}_{i=1}^{3}$ from 8, 16, and 24 layers of the frozen CLIP visual encoder. These features are then fused via a FPN module Wang et al. (2022) and decode through a Transformer block to generate preliminary multi-modal features $F_{m}$. To strengthen the alignment between text and pixels and emphasize class semantics of object, we introduce a cross-modal multi-head attention (CMA). The computation process follows the basic attention mechanism, projecting features from both modalities into different embedding spaces, and calculating the correlation between text and image features across multiple subspaces. This highlights the class semantic knowledge of object in the image, ultimately generating the feature $F_{c}$. Furthermore, considering the effectiveness of multi-level feature interactions in image understanding, we propose a Multi-level Visual Collaboration module (MVCM) to enrich visual features at different levels and generate the feature $F_{v}$. Fig. 4 visualizes the stages within the CPG, demonstrating progressive improvements in object localization and validating the effectiveness of our approach.
Multi-level Visual Collaboration Module (MVCM). Existing research has shown that the collaboration between low and high level structural information is essential for fine-grained segmentation, reflecting principles akin to human visual processing. In order to take advantage of the coarse segmentation multi-level features of the CLIP visual encoder to generate more fine-grained visual features, we propose the MVCM to enhance the multi-layer features $F_{c}$, $F_{2}$ and $F_{3}$. Specially, this module consists of two stages: the alignment stage and the enhancement stage. In the alignment stage, we employ two Multi-Head Self-Attention (MHSA) modules to align $F_{c}$ with $F_{2}$ and $F_{c}$ with $F_{3}$, followed by a summation operation to fuse them, generating intermediate features $F_{\text{n}}$. These features are then further enhanced.
$F_{\text{n}}=\text{MHSA}(F_{c},F_{2})\oplus\text{MHSA}(F_{c},F_{3})$ (1)
In the enhancement stage, we apply a residual-based progressive refinement strategy on $F_{\text{n}}$. This strategy progressively strengthens the feature, ultimately producing the fused feature representation $F_{v}$. This design aims to maximize the retention of fine details while extracting rich semantic information, thereby improving the model’s ability to detect camouflaged objects. which are formulated as:
$F_{n}^{1}=\text{Conv}(\text{Conv}(F_{n})\oplus F_{n})$ (2)
$F_{n}^{2}=\text{Conv}(\text{Conv}(F_{n}\otimes F_{n}^{1})\oplus F_{n}\otimes F_{n}^{1})$ (3)
$F_{n}^{3}=\text{Conv}(\text{Conv}(F_{n}\otimes F_{n}^{2})\oplus F_{n}\otimes F_{n}^{2})$ (4)
$F_{v}=\text{Conv}(\text{Concat}(F_{n}^{1},F_{n}^{2},F_{n}^{3})).$ (5)
Here, Conv() signifies a $3\times 3$ convolution.
Figure 4: Visualization of the internal feature in the CPG.
4.2 Class-Guided Detector (CGD)
Overview. To improve the performance of COD, we further propose a novel detector CGD. Which leverages class-promt information to accurately segment camouflaged objects and suppress background noise. Given an input $I_{r}$ of size $W\times H$, We first use PVTv2 Wang et al. (2021) as the visual backbone to extract a set of depth features $\{X_{i}\}_{i=0}^{4}$. As shown in Fig. 3, the semantic-rich feature $f_{4}$, the feature $F_{v}$, and the initially fused multi-modal information $F_{m}$ are input into the class-semantic guidance (CSG). The module aims to explore the internal correlations between the backbone features and the other modalities from both shallow and deep layers, thus generating class-semantic guidance map $G_{c}$ that embed both class semantics and camouflaged semantics. These features $\{X_{i}\}_{i=1}^{4}$ and $G_{c}$ are then passed to the semantics consistency module to further extract fine-grained features related to the camouflaged object, ultimately leading to an accurate segmentation. The process iteratively accumulates intermediate segmentation results to decouple the final output $\{P_{i}\}_{i=1}^{4}$.
Semantics Consistency Module (SCM). Traditional COD models rely on a visual backbone that extracts camouflaged semantics containing both foreground and background clues, often resulting in false negatives and false positives. To address these issues, we propose the SCM. By class guided features, the SCM pays more attention to features with high correlation with guided features, alleviating the interference of irrelevant noise information. Most camouflage is achieved by using channel information (such as color) and spatial location in relatively few cases. So the SCM consists of two stages: the first stage, spatial localization, focuses on capturing spatial information of visual features related to guidance feature, while the second stage, channel refinement, explores inter-channel relationships and adaptively selects features that align with class semantics to suppress background interference. Specifically, we utilize two parallel branches to process multi-level features $X_{i}$ and $X_{i+1}$ independently, facilitating their interaction with class semantic knowledge. Each branch incorporates weighted summation and dot product operations, ensuring a comprehensive interaction between the object class semantic features and the multi-level features. This interaction establishes an initial alignment between the camouflaged foreground and background semantics and the class-guided semantic features. Consequently, it produces a new class-semantic relevance matrix $G^{\prime}_{\text{c}}$ and enriched class semantic features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$. The process is formulated as follows:
$R_{\text{cls}}=\text{Softmax}\left(\text{Conv}\big{(}X_{i}\oplus X_{i+1}\big{)}\otimes G_{\text{c}}\right)$ (6)
$\hat{X}_{i}=R_{\text{cls}}\otimes X_{i},\quad\hat{X}_{i+1}=R_{\text{cls}}\otimes X_{i+1}$ (7)
$R^{\prime}_{\text{cls}}=\text{FC}\left(\text{Concat}\big{(}\hat{X}_{i+1},\hat{X}_{i}\big{)}\right)$ (8)
$\tilde{X}_{i}=R^{\prime}_{\text{cls}}\oplus\hat{X}_{i},\quad\tilde{X}_{i+1}=R^{\prime}_{\text{cls}}\oplus\hat{X}_{i+1}$ (9)
Here, $i$ denotes the $i$-th operator, $\text{FC}()$ refers to linear projection, while $\text{Conv}()$ signifies a $3\times 3$ convolution.
Table 1: Comparison with recent COD and RIS methods with different training settings. “+”: optimizing the detector under our class prompt generatoron the proposed dataset CamoClass; “-”: methods without publicly available results; CGD: the proposed class-guided detector; 2D Prompt: domain-specific prompts Luo et al. (2024); COD-TAX: a coarse-grained text set generated by GPT-4V for camouflaged environments Zhang et al. (2024). The best three results are highlighted in red, green, and blue fonts.
Methods Prompt Backbone CAMO (250 images) COD10K (2026 images) NC4K (4121 images) CHAMELEON (76 images)
$S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{\phi}^{m}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{\phi}^{m}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{\phi}^{m}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{\phi}^{m}\uparrow$
Referring Image Segmentation
CRIS Wang et al. (2022) CamoClass CLIP 0.305 0.132 0.457 0.366 0.258 0.082 0.583 0.281 0.278 0.129 0.533 0.311 0.251 0.117 0.581 0.273
CGFormer Tang et al. (2023) CamoClass BETR, Swin 0.285 0.177 0.606 0.285 0.273 0.086 0.632 0.262 0.279 0.147 0.617 0.276 0.293 0.133 0.597 0.286
ASDA Yue et al. (2024) CamoClass CLIP 0.790 0.643 0.091 0.819 0.807 0.601 0.048 0.863 0.847 0.701 0.059 0.878 0.832 0.647 0.063 0.864
Camouflaged Object Detection
SINet Fan et al. (2020a) None ResNet 0.751 0.606 0.100 0.771 0.751 0.551 0.051 0.806 0.808 0.723 0.058 0.887 0.869 0.740 0.044 0.899
PFNet Mei et al. (2021) None ResNet 0.782 0.695 0.085 0.855 0.800 0.660 0.040 0.877 0.829 0.745 0.053 0.887 0.882 0.810 0.033 0.942
BGNet Sun et al. (2022) None Res2Net 0.812 0.749 0.073 0.870 0.831 0.722 0.033 0.901 0.851 0.788 0.044 0.907 0.901 0.851 0.027 0.943
ZoomNet Pang et al. (2022) None ResNet 0.820 0.752 0.066 0.877 0.838 0.729 0.029 0.888 0.853 0.784 0.043 0.896 0.902 0.845 0.023 0.952
FEDER He et al. (2023) None Res2Net 0.836 0.807 0.066 0.897 0.844 0.748 0.029 0.911 0.862 0.824 0.042 0.913 0.887 0.834 0.030 0.946
FSPNet Huang et al. (2023) None VIT 0.851 0.802 0.056 0.905 0.850 0.755 0.028 0.912 0.879 0.816 0.035 0.914 0.908 0.868 0.022 0.943
HitNet Hu et al. (2022) None PVT 0.849 0.809 0.055 0.906 0.871 0.806 0.023 0.935 0.875 0.834 0.037 0.926 0.921 0.897 0.019 0.967
PRNet Hu et al. (2024) None PVT 0.872 0.831 0.050 0.922 0.874 0.799 0.023 0.937 0.891 0.848 0.031 0.935 0.914 0.874 0.020 0.971
CamoFormer Yin et al. (2024) None PVT 0.872 0.831 0.046 0.929 0.869 0.786 0.023 0.932 0.892 0.847 0.030 0.939 0.910 0.865 0.022 0.957
FSEL Sun et al. (2024) None PVT 0.885 0.857 0.040 0.942 0.877 0.799 0.021 0.937 0.892 0.852 0.030 0.941 - - - -
RISNet Wang et al. (2024) Depth Map PVT 0.870 0.827 0.050 0.922 0.873 0.799 0.025 0.931 0.882 0.834 0.037 0.926 - - - -
VSCodes Luo et al. (2024) 2D Prompt Swin 0.873 0.820 0.046 0.925 0.869 0.780 0.025 0.931 0.882 0.841 0.032 0.935 - - - -
DSAM Yu et al. (2024) Depth Map SAM, PVT 0.832 0.794 0.061 0.913 0.846 0.760 0.033 0.931 0.871 0.826 0.040 0.932 - - - -
ACUMEN Zhang et al. (2024) COD-TAX CLIP 0.886 0.850 0.039 0.939 0.852 0.761 0.026 0.930 0.874 0.826 0.036 0.932 - - - -
SINet+ Fan et al. (2020a) CamoClass CLIP, ResNet 0.870 0.807 0.047 0.920 0.857 0.733 0.029 0.912 0.882 0.806 0.038 0.925 0.882 0.795 0.036 0.927
PFNet+ Mei et al. (2021) CamoClass CLIP, ResNet 0.873 0.819 0.045 0.925 0.860 0.742 0.028 0.919 0.890 0.817 0.035 0.925 0.889 0.816 0.031 0.947
BGNet+ Sun et al. (2022) CamoClass CLIP, Res2Net 0.879 0.836 0.042 0.929 0.875 0.788 0.022 0.936 0.893 0.845 0.030 0.939 0.911 0.858 0.025 0.951
FSPNet+ Huang et al. (2023) CamoClass CLIP, VIT 0.879 0.832 0.041 0.939 0.878 0.768 0.025 0.933 0.881 0.830 0.032 0.940 0.911 0.870 0.022 0.956
GGD (Ours) None PVT 0.876 0.838 0.045 0.932 0.872 0.793 0.023 0.935 0.889 0.847 0.032 0.935 0.917 0.876 0.021 0.965
CGNet (Ours) CamoClass CLIP, PVT 0.894 0.863 0.036 0.946 0.890 0.824 0.018 0.948 0.903 0.868 0.027 0.948 0.931 0.902 0.017 0.972
To enhance the feature expression and focus on class-relevant information, we propose a channel-adaptive refinement mechanism. Initially, features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$ are concatenated along the channel dimension to form a unified feature representation $X_{\text{cot}}$. This representation is then partitioned into four subgroups $\{g_{1},g_{2},g_{3},g_{4}\}$ along the channel dimension. For each subgroup $g_{i}$, its concern weight $w_{i}$ is calculated through the linear layer and sigmoid activation function. These attention weights modulate the importance of each subgroup by element-wise multiplication, emphasizing features that are consistent with the semantics of the object class. The weighted subgroups are subsequently fused along the channel dimension to produce the refined feature representation $F_{\text{s}}$, which highlights the critical information associated with the target class. The process is defined as follows:
$X_{\text{cot}}=\text{Conv}(\text{Concat}(\tilde{X}_{i},\tilde{X}_{i+1}))$ (10)
$\{g_{1},g_{2},\dots,g_{n}\}=\text{Split}(X_{\text{cot}})$ (11)
$w_{i}=\text{Gate}(g_{i})$ (12)
$g^{\prime}_{i}=\text{Conv}_{i}(x_{\text{cat}}\otimes w_{i})$ (13)
$F^{i}_{\text{s}}=\text{Concat}(g^{\prime}_{1},g^{\prime}_{2},\dots,g^{\prime}_{n})$ (14)
Here, Gate()()( ) refers to tow linear projection and a sigmoid, while Conv()()( ) signifies a $3\times 3$ convolution.
4.3 Loss Function
Since our ultimate goal is to generate a binary foreground graph, a structured loss function consisting of BCE loss and IoU loss can be used in the optimization process Wei et al. (2020). Specifically, the loss function can be expressed as:
$\mathcal{L}(P,G)=\sum_{i=1}^{6}\mathcal{L}_{\text{bce}}(P_{i},G)+\mathcal{L}_{\text{iou}}(P_{i},G),$ (15)
where $P=\{F_{m},F_{v},P_{4},P_{3},P_{2},P_{1}\}$ represent the predictions generated by our model, and $G$ refers to the ground truth annotation.
5 Experiments
5.1 Implementation Details
Data Settings. We used the CamoClass dataset to evaluate our approach. The training set consists of CAMO Le et al. (2019) and COD10K Fan et al. (2022), with a total of 4040 samples and their class-level textual information. The test set consists of test samples from four benchmark datasets, totaling 8615 samples and their class-level textual information. It is important to note that we have fine-grained segmentation of the multi-class samples in the benchmark dataset, so that our total number of samples exceeds the original number of these four benchmark datasets.
Evaluation Metrics. In this work, we use five widely recognized and standard metrics to assess the performance of our method, including S-measure ($S_{\alpha}$an et al. (2017), E-measure ($E^{m}_{\phi}$an et al. (2018), weighted F-measure ($F_{\beta}^{\omega}$v et al. (2021), average F-measure ($F^{m}_{\beta}$v et al. (2021) and mean absolute error ($M$Perazzi et al. (2012).
Training Settings. Following the setup in previous studies Xu et al. (2023, 2022), the parameters of the CLIP model were frozen during the training process, while the parameters of the remaining network components were learnable. We use the Adam optimizer with as learning rate of $1\times 10^{-4}$ to optimize the network parameters, with a batch size of 10 and a total of 200 epochs. We apply various data augmentation techniques to the training data, including random cropping, horizontal flipping, and color jittering. Specially, during both the training and inference phases, the image input size for the CLIP model was $336\times 336$ pixels, while the input size for the class-guided detector was $448\times 448$ pixels. All experiments are conducted on a server equipped with an NVIDIA RTX 3090 (with 24GB memory).
Figure 5: Visual comparisons of prediction maps produced by different COD approaches.
5.2 Comparison with State-of-the-Art Methods
To assess our approach, we compare CGNet against state-of-the-art COD and RIS methods on four benchmark datasets, as summarized in Tab. 1. We also retrain certain COD and RIS methods on the CamoClass dataset, utilizing the CPG and CSG to gennerate prompt feature as $X_{4}$ input to enhance their performance.
Quantitative Evaluation. As shown in Tab. 1, CGNet consistently surpasses competing methods across all datasets, showing a significant and substantial performance advantage. And CGNet are on average 3.5% and 5.5% better than the second best method. Due to CRIS, CGFormer, and ASDA are not specifically designed for the COD task, their performance on the four datasets is poor. Moreover, we also present the results of detectors optimized under the class prompt generator (CPG) and proposed CamoClass. In Tab. 1, CPG improve other detectors by 19.73% (SINet), 13.3% (BGNet), 12.68% (PFNet), 4.43% (FPSNet), and increase our class-guided detector by 6.24% (PVT), verifying that CPG serve as effective plug-and-play module for performance improvement.
Table 2: Comparison of methods on seen and unseen classes ($\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$), the best results are marked in red.
Model Seen (2212 images) Unseen (4261 images)
$S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F^{m}_{\beta}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F^{m}_{\beta}\uparrow$
SINet 0.778 0.636 0.043 0.684 0.804 0.718 0.060 0.765
PFNet 0.802 0.666 0.040 0.707 0.826 0.742 0.055 0.781
BGNet 0.833 0.728 0.033 0.759 0.848 0.785 0.046 0.817
FEDER 0.824 0.721 0.032 0.756 0.844 0.785 0.046 0.821
ZoomNet 0.840 0.733 0.030 0.770 0.851 0.782 0.045 0.815
HitNet 0.873 0.810 0.024 0.826 0.873 0.831 0.038 0.851
PRNet 0.875 0.804 0.022 0.826 0.890 0.846 0.032 0.868
CGD 0.874 0.799 0.023 0.819 0.891 0.848 0.032 0.868
CGNet 0.892 0.828 0.019 0.845 0.902 0.863 0.028 0.880
Table 3: Ablation comparison of proposed components. C: semantics consistency module; D: CLIP; E: class-semantic guidance; F: cross-modal multi-head attention; G: multi-level visual collaboration module.
No. Model Param. MACs FPS COD10K (2026 images)
(M) (G) $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $M\downarrow$
I Baseline 24.85 7.80 78.91 0.824 0.914 0.792 0.041
II +C 29.08 17.42 44.40 0.872 0.935 0.793 0.023
III +C,D 321.49 78.13 14.18 0.875 0.924 0.791 0.022
IV +C,D,E 327.01 87.34 12.89 0.880 0.928 0.803 0.021
V +C,D,E,F 342.38 88.94 12.68 0.889 0.937 0.818 0.019
VI +C,D,E,G 346.78 90.87 12.48 0.885 0.934 0.809 0.020
VII +C,D,E,F,G 371.33 101.04 11.20 0.890 0.948 0.824 0.018
Qualitative Evaluation. Fig. 5 presents a visual comparison between our proposed CGNet method and other advanced methods. We evaluated challenging camouflaged images featuring different object types under various extreme scenes. These scenes include multiple objects (the first line), structural degradation (second row), occluded scenes (third row), edge similarity (fourth row), and small objects (last row). Such conditions can confound existing COD methods, leading to localization errors, blurred edges, and incomplete segmentation. As observed in Fig. 5, in the first line, other advanced methods struggle to completely segment the two owls due to the high internal similarity of the samples. In second row, the structural semantics of the katydid are disrupted, causing confusion in segmentation for other methods. In third row, heavy occlusion results in interrupted semantic information, preventing other methods from fully segmenting the target. Similarly, In forth and fifth row, complex edge details and small object sizes challenge other methods, resulting in coarse or incomplete segmentation. In contrast to the aforementioned methods, our CGNet approach overcomes these challenges and demonstrates superior predictive performance across these extreme scenes.
5.3 Ablation Study
Impact of Components. To analyze the effect of different components in detail, we evaluate their performance in Tab. 3. In addition, Fig. 6 shows the predicted results for each component. The groups in the table, from I to VII, are incrementally added with one module at a time, where I represents the base network. Our analysis demonstrates that excluding any component leads to a degradation in the performance of the CGNet model. II represents our proposed the class-guided detector. Tab. 3 and Fig. 6 show that CGD achieves excellent performance on COD10K. However, it fails to completely segment targets with a high degree of camouflage. IV performs better than II and III, which indicates the ability of class-Semantic guidance to mitigate domain differences from multiple sources of visual features. However, it still exhibits some false positives and false negatives in complex environments. Compared to I (Baseline), VII (CGNet) improves the performance by 8.3%, which indicates the model progressively focuses on features that are highly relevant to the object class semantics and achieves complete object segmentation under the guidance of class-level textual information. Importantly, incorporating the “D” component significantly increases the model’s parameter count, mainly due to the addition of the CLIP encoder.
Figure 6: Visual results of the effectiveness of our components, the C, D, E, F, and G have already been explained in Tab. 3.
Performance on Unseen and Seen Classes. As stated in Sec. 3, we categorize the CamoClass test samples into two sets $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$. This division allows us to evaluate the generalization capability of different methods. Since the distribution of $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$ varies significantly, we rely on results provided by the authors for a fair comparison. As shown in Tab. 2, we evaluate the performance of various COD methods on them. Compared to other COD methods, the results demonstrate that our CGNet achieves superior performance on both sets, particularly on $\mathcal{C}_{unseen}$. This indicates the strong generalization ability of our framework, which benefits from the integration of class-level textual information. To more intuitively show how each method performs on $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$, we provide a comparison of visual results in the supplementary material.
6 Conclusion
In this paper, we propose a novel task, termed Class-Guided Camouflaged Object Detection, which employs class-level textual information to guide COD. To facilitate this task, we have meticulously constructed a large-scale dataset, CamoClass, which not only comprises real-world images under diverse and complex conditions, but also includes corresponding class text of objects. Building upon this foundation, we introduce a multi-stage reasoning framework, which include the class prompt generator (CPG) and the class-guided detector (CGD). Extensive experiments demonstrate the effectiveness of the proposed method and its superior performance compared with the existing COD method. Extensive experiments prove that our method has strong scalability in COD and can successfully detect objects of the unseen classes. Moreover, as a plug-and-play component, the CPG significantly enhances the performance of existing COD detectors. We hope that our exploration will inspire future research, for instance, investigating multi-object camouflage segmentation and examining whether COD models truly capture the semantics of camouflage.